### v1

# Obfuscation-Resilient Binary Code Similarity Analysis using Dominance Enhanced Semantic Graph

Abstract.
Binary code similarity analysis (BCSA) serves as a core technique for binary analysis tasks such as vulnerability detection. While current graph-based BCSA approaches capture substantial semantics and show strong performance, their performance suffers under code obfuscation due to the unstable control flow. To address this issue, we develop ORCAS, an Obfuscation-Resilient BCSA model based on Dominance Enhanced Semantic Graph (DESG). The DESG is an original binary code representation, capturing more binaries’ implicit semantics without control flow structure, including inter-instruction relations, inter-basic block relations, and instruction-basic block relations. ORCAS robustly scores semantic similarity across binary functions from different obfuscation options, optimization levels, and instruction set architectures. Extensive evaluation on the BinKit dataset shows ORCAS significantly outperforms eight baselines, achieving an average 12.1% PR-AUC gain when using combined three obfuscation options compared to the state-of-the-art approaches. Furthermore, ORCAS improves recall by up to 43% on an original obfuscated real-world vulnerability dataset, which we released to facilitate future research.
Binary Code Similarity Analysis, Obfuscation-Resilient, Dominator Tree
ccs: Security and privacy Software reverse engineeringccs: Computing methodologies Machine learning
1. Introduction
Code obfuscation obscures the code’s control flow and basic blocks while preserving the function semantics, which has been widely applied to increase the difficulty of unauthorized reverse engineering (Gao et al., 2023), e.g., in 2018, roughly 50% of popular apps on Google Play with over 10 million downloads were obfuscated to protect against plagiarism and repackaging (Wermke et al., 2018). While malware authors also exploit code obfuscation techniques to conceal malicious contents and evade detection (Cheng et al., 2021; Cuiying et al., 2024). In 2024, approximately eight million new malware samples are identified, with over 300,000 being detected each day. [^1] https://portal.av-atlas.org/malware
Binary code similarity analysis (BCSA) scores the semantic similarity between two given binary code snippets, which is commonly used for malware identification (Cesare et al., 2013; Kim et al., 2019), vulnerability detection (Luo et al., 2023; Li et al., 2023), patch analysis (Hu et al., 2016; Xu et al., 2017a; Kargén and Shahmehri, 2017), and software plagiarism detection (Luo et al., 2017), etc. Existing BCSA approaches perform well in BCSA tasks across different compilers, optimization levels, and instruction set architectures (ISAs). However, they face challenges when analyzing obfuscated binary code. To be specific, a binary function can be represented as a control flow graph (CFG), where a node represents a basic block, and the directed edge between two nodes represents their control flow. There are three widely used code obfuscation techniques: instruction substitution (SUB), bogus control flow (BCF), and control flow flattening (FLA) (Junod et al., 2015). SUB obfuscation replaces simple instructions with semantically equivalent complex instructions, with control flow untouched. BCF obfuscation modifies the control flow by adding bogus basic blocks and edges. While FLA obfuscation complicates the control flow with a complex hierarchy, and then flattens the control flow into a linear form. Meanwhile, different techniques can be combined to obfuscate the code, e.g., for a given CFG, first, BCF is applied to double the number of nodes and edges, then FLA is applied to double them again (Ding et al., 2019). Then we can see that a CFG and its obfuscated CFG can have different nodes, different flow structures, and even arbitrarily added bogus nodes and edges, which present great challenges to existing BCSA approaches to detect their semantic similarity.
According to how the binary code is represented, existing BCSA approaches can be categorized into instruction stream based and graph based. When instruction stream based approaches are used, binary code is first translated into assembly code or intermediate representation (IR) code, which are treated as natural language sentences. Then natural language processing (NLP) technologies are applied to capture the code’s implicit semantics. NLP approaches have strong contextual understanding capabilities, however, they usually rely on pre-training tasks tailored to binary code, which require large-scale datasets and expensive computational resources (Wang et al., 2022; Li et al., 2021; Ahn et al., 2022; Yu et al., 2020; Luo et al., 2023). In addition, obfuscating simple instructions with complex ones increases contextual complexity and token length, which degrades the performance of NLP technologies.
Figure 1. The D-tree and PD-tree of the CFG before and after the application of the BCF and FLA obfuscation, where gray nodes denote fake basic blocks generated by the obfuscation techniques.
When graph based approaches are applied, a binary function is represented as a feature enhanced CFG or semantics-oriented graph (SOG). Then graph neural network (GNN) technologies are used to learn the code’s semantics. First, various features are incorporated into the CFG for capturing code’s semantics: (1) Incorporating basic block features (e.g., number of instructions), which are obtained by manually crafting (Xu et al., 2017b; Feng et al., 2016) or automatically generating using NLP technologies (Luo et al., 2023; Yu et al., 2020; Jia et al., 2024a); (2) Incorporating data flow graph (DFG) and call graph (CG) to provide richer semantic information between instructions and functions (Zhang et al., 2024; Duan et al., 2020; Guo et al., 2022). Second, when a binary function is represented as a SOG (He et al., 2024), nodes represent instruction opcodes and operands and edges reveal intra-instruction as well as inter-instruction relations, where control flow between basic blocks are represented using inter-instruction relations. Graph based approaches capture more binary code semantics and demonstrate remarkable BCSA performance. However, when code obfuscation technologies are applied, the unstable control flow structure degrades graph based BCSA performance (Ding et al., 2019), this problem remains unresolved.
This paper makes 4 main contributions for obtaining obfuscation-resilient BCSA:
  • Propose an original Dominance Enhanced Semantic Graph (DESG) embedding for representing a binary function, removing the control flow structure while revealing inter-basic block relations using dominance and post-dominance relations with a constructed dominator tree, and meanwhile revealing other graph semantics of existing BCSA approaches.
  • Develop ORCAS, a novel Obfuscation-Resilient BCSA model, capturing more binaries’ implicit semantics without control flow structure, for more robust BCSA.
  • Construct an original obfuscated real-world vulnerability dataset to evaluation of 1-day vulnerability detection on different obfuscation techniques.
  • Conduct extensive experiments to evaluate the performance of ORCAS against baselines over the existing dataset and our constructed dataset, and experimental results demonstrate the effectiveness of ORCAS.
The rest of the paper is organized as follows. Section 2 introduces our motivation. Section 3 presents our original BCSA model ORCAS. Section 4 describes the experimental setup. Section 5 describes the performance evaluation of ORCAS. Section 6 reviews the related work and finally Section 7 concludes the paper.
2. Motivation
This section describes the motivation behind this paper. In general, BCSA can be conducted across different granularity, ranging from coarse to fine: program-level, function-level, slice-level, basic block-level, and instruction-level (Haq and Caballero, 2021). This paper focus on the function-level BCSA since functions are considered basic building blocks of program functionality, function-level BCSAs (Wang et al., 2022; Xu et al., 2023; Luo et al., 2023; He et al., 2024; Zou et al., 2022; Ding et al., 2019; Pei et al., 2023; Feng et al., 2024) enables many real-world applications such as patch analysis (Feng et al., 2019; Yang et al., 2023b), and can be easily extended to detect similarities between whole programs.
Definition 2.1 (Function).
A function is an ordered sequence of binary instructions compiled from a function written in high-level programming languages, based on different settings, e.g., obfuscation options, optimization levels, and ISAs. A binary function possesses specific semantics and structures, which can be represented as assembly code, IR code, a graph (e.g., CFG), or a tree (e.g., abstract syntax tree).
Actually, a binary function can be represented as a CFG, denoted as $G=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}=\{v_{0},v_{1},\ldots,v_{n-1}\}$ is the set of nodes, $v_{0}$ and $v_{n-1}$ denotes the unique entry and exit node respectively. =$\mathcal{E}=\{(v_{i},v_{j})\mid v_{i},v_{j}\in\mathcal{V}\}$ is the set of directed edges between nodes. To be specific, a node represents a basic block, which is a sequence of consecutive instructions ending with a branch or return instruction.
To put the discussion into perspective, an example CFG $G$ is depicted in Fig. LABEL:fig:motivation(a) to illustrate the problem. SUB obfuscates instructions, i.e., the nodes in a CFG, while FLA and BCF obfuscate nodes as well as edges in a CFG, which are widely adopted by various BCSA approaches (He et al., 2024; Li et al., 2021; Guo et al., 2022; David et al., 2017; Jia et al., 2024b; Luo et al., 2023) for more notable performances. Then, BCF and FLA obfuscation have been applied to CFG $G$ and obtained an obfuscated CFG $\tilde{G}$, which is depicted in Fig. LABEL:fig:motivation(d), where the dashed boxes display a typical application of an official Obfuscator-LLVM’s BCF and FLA obfuscation techniques (Junod et al., 2015). We can see that node $v_{3}$ and node $v_{4}$ have an edge to node $v_{5}$ in CFG $G$, while in the obfuscated CFG $\tilde{G}$, neither node $v_{3}$ nor node $v_{4}$ has an edge to node $v_{5}$. That is, the control flow structure of CFG $G$ is different from that of CFG $\tilde{G}$.
From a CFG, a dominator tree (D-Tree) and a post-dominator tree (PD-Tree) can be constructed to capture the dominance and post-dominance relations respectively. Figs. LABEL:fig:motivation(b) and LABEL:fig:motivation(c) depicts the D-Tree and PD-Tree constructed upon CFG $G$. Generally, in a D-Tree, node $v_{i}$ dominates node $v_{j}$ if and only if node $v_{i}$ is included in every path from the entry node $v_{0}$ to node $v_{j}$, in this case node $v_{i}$ is the dominator of node $v_{j}$. Similarly, in a PD-Tree, node $v_{j}$ post-dominates node $v_{i}$ if and only if node $v_{j}$ is included in every path from the node $v_{i}$ to the exit node $v_{n-1}$, in this case node $v_{j}$ is the post-dominator of node $v_{i}$. The dominance relations help identify the key nodes from the entry node to a specific node, while the post-dominance relations help identify the key nodes from a specific node to the exit node.
Table 1. Average GED across different sizes of basic blocks.
Basic Block Size Range (0,50] (50,100] (100,150] ¿150
CFG 207 553 916 1084
D-Tree 162 431 734 828
Figs. LABEL:fig:motivation(e) and LABEL:fig:motivation(f) depict $\tilde{G}$’s D-Tree and PD-Tree respectively. We can see that there exists a dominance relation where node $v_{0}$ dominates nodes $v_{6},v_{2},v_{1}$, and node $v_{2}$ dominates nodes $v_{3},v_{4},v_{5}$ in the D-Tree of CFG $G$, and such dominance relations for node $v_{0}$ and node $v_{2}$ remain unchanged in the D-Tree of CFG $\tilde{G}$ in Fig. LABEL:fig:motivation(e). Similarly, there are post-dominance relations from node $v_{6}$ to nodes $v_{1},v_{5},v_{0}$, and from node $v_{5}$ to nodes $v_{2},v_{3},v_{4}$ in the PD-Tree of CFG $G$, and we can also have such post-dominance relations retain in the PD-Tree of CFG $\tilde{G}$ in Fig. LABEL:fig:motivation(f).
Based on the aforementioned discussion, we observe that dominance and post-dominance relations exhibit superior structural stability compared to control flow structure under obfuscation. To quantify this stability difference, we employ graph edit distance (GED), similar to (Pawlowski et al., 2016), which measures the minimum number of edit operations (i.e., node/edge additions/deletions or node relabelling) required to transform one graph into another. Specifically, we compute GED between original and obfuscated versions of both CFGs and D-Trees using approximately 1,800 function samples obfuscated with combined BCF and FLA obfuscation. Table 1 presents a comparative analysis of structural stability between CFG and D-Tree under obfuscation across sizes of basic blocks. Notably, D-Trees exhibit a 23.6% lower average GED for functions with basic block sizes exceeding 150. This demonstrates that D-Tree consistently exhibit superior structural stability compared to CFG under obfuscation.
3. ORCAS
This section elaborates on how ORCAS is designed. Fig. 2 depicts the ORCAS model, which takes two binary functions as input and decompiles them using Ghidra (Eagle and Nance, 2020) into P-Code IR functions. In the function embedding phase, we construct the DESGs and employ a Gated Graph Neural Network (GGNN) (Li et al., 2015) to generate embeddings. In the similarity computation phase, we utilize cosine similarity to measure binary function pairs’ similarity using their embeddings.
Figure 2. The framework of ORCAS. It first translates binary functions into P-Code functions, generates function embeddings, and then computes their cosine similarity.
3.1. Intermediate Representation
To mitigate the differences between various ISAs and simplify instruction types, IRs are widely utilized to represent binary functions in recent BCSA approaches (Yu et al., 2023; He et al., 2024; Luo et al., 2023). Since Ghidra outperforms other decompilers like IDA-Pro (Eagle, 2011) and Angr (Shoshitaishvili et al., 2016) in correctly identified function starts task (Shaila et al., 2021), we utilize Ghidra to decompile the binary code and lift it into P-Code, Ghidra’s IR.
According to reference manual [^2] https://spinsel.dev/assets/2020-06-17-ghidra-brainfuck-processor-1/ghidra_docs/language_spec/html/pcoderef.html, P-Code is a register transfer language designed for reverse engineering applications. It translates processor instructions into a sequence of P-Code operations, each P-Code operation takes one or more varnodes as input and may produce a single output varnode. A varnode is a generalized abstraction of a register or memory location, represented by a formal triple: (address space type, offset, size), where the offset denotes the displacement within a specific type of address space, and the size defines the allocated byte length.
The address space for P-Code is a generalization of random access memory and is defined simply as an indexed sequence of bytes that can be read and written by the P-Code operations. There are five types of address spaces: ram, register, constant, unique, and stack. Ram address space represents an abstraction of the physical memory accessed by the processor, while register address space refers to the processor’s general-purpose registers. Unique address space holds intermediate value generated during processing, and stack address space denotes the memory space accessed through the stack pointer. Finally, constant address space is reserved for storing fixed values.
Table 2. Varnode normalization strategies.
Address Space Before Normalization After Normalization
Ram (ram,0x8,8) ram or abs
Register (register,0x0,4) x86_r_0 or ARM_r_0
Constant (const,0x0,4) c_0
Unique (unique,0x0,4) unique
Stack (stack,0xf,8) stack
Varnode Normalization. Since each binary function contains too many different varnodes (e.g., with various offsets), it is difficult for the GNN’s vocabulary to encompass all varnodes. To tackle this issue, we propose varnode normalization strategies, as outlined in Table 2. We first disregard the size of all varnodes and then apply tailored strategies for different address spaces to reduce the vocabulary size:
  • For ram address space, if the varnode represents a standard library function, we normalize it using the function name; if not, we normalize it with the identifier ram. Notably, standard library function retain their original function name by default under obfuscation, as symbol table-based dynamic linking would fail with obfuscated name (Picheta, 2020).
  • For register address space, due to the varying number of registers across ISAs, we retain the offset while incorporating an architecture identifier (i.e., ARM or x86).
  • For constant address space, we retain the offset and normalize the varnode with the identifier c_offset.
  • For unique and stack address spaces, we normalize them to the identifiers unique and stack, respectively.
3.2. DESG Construction
To capture the comprehensive and stable semantics of the obfuscated binary function, we propose a graph based representation called the DESG. We first discuss the implicit semantics of the binary function, and then elaborate on how the DESG is constructed.
3.2.1. Implicit Semantics of Binary Function
Different to natural language, binary code has well-defined implicit semantics, which can be categorized into inter-basic block relations, inter-instruction relations, and instruction-basic block relations.
Inter-basic block relations. Control flow relations define the original relations between basic blocks in a CFG. However, bogus control flow generated by obfuscation techniques may introduce misleading information. As a result, we chose to remove them. Based on CFG, we introduce two new relations between basic blocks: dominance and post-dominance relations, derived from dominance analysis (Lengauer and Tarjan, 1979; Azen and Budescu, 2003). These relations are crucial for understanding control flow structure and analyzing the execution paths of obfuscated binary function.
Inter-instruction relations. We further divide inter-instruction relations into two categories: data relation and effect relation. In previous binary analysis works (Li et al., 2021; David et al., 2017; He et al., 2024; Guo et al., 2022; Jia et al., 2024b; Luo et al., 2023), the def-use relation between instructions is often used since it is a fundamental concept in program analysis. We adopt the def-use relation as the data relation, which reveals how instruction use values are defined by others. Inspired by (Click and Paleczny, 1995; He et al., 2024), effect relation imposes restrictions on the execution order between instructions, which is derived from def-use relations and it is essentially a special case of the def-use relation. Especially the effect relation focuses on the execution order between a CALL instruction and an instruction that may read from or write to memory. Take Fig. 3a as an example, the instruction r2=abs(r1) may use the value r1 defined by the memory read instruction r1=LOAD(0x8). This establishes an effect relation between them, indicating that it must be ensured that the instruction r1=LOAD(0x8) executes before the instruction r2=abs(r1).
Instruction-basic block relations. A CFG inherently implies that instructions are contained within specific basic blocks. To explicitly expose this to the GNN, we introduce a contain relation between basic blocks and instructions, indicating that an instruction belongs to a particular basic block. This relation directly reduces the distance between instructions within the same basic block in the graph, facilitating information interaction during the GNN training phase.
Figure 3. An example of constructing the DESG. For clarity, each basic block in (b) shows only the core logic P-Code operations.
3.2.2. Steps to the DESG
Fig. 3 presents the construction process of the DESG, which can intuitively be divided into five steps. The details are outlined in Algorithm 1.
Step ❶. We first decompile the binary function to obtain the CFG in the original P-Code, and then apply varnode normalization strategies to generate the CFG in the normalized P-Code function in Fig. 3b.
Step ❷. We split the instruction sequence in each basic block into individual instructions, where a node represents an individual instruction, while an edge represents the relation between instructions (i.e., data relation or effect relation). Thus, we obtain the instruction-based semantic graph (ISG).
Step ❸. Based on the ISG, we further reveal the structure within all instructions by splitting the opcode and operand(s) (i.e., varnodes) into tokens. As a result, the relations between instructions are refined into relations between tokens, with the effect relation specifically defined on the opcode tokens. Similar to (He et al., 2024), the relation between opcode tokens and operand tokens can be considered a data relation, which we refine into two distinct relations: one for the first operand and one for the second operand (if exists). For example, in the instruction x86_r_0 = INT_SUB x86_r_0, c_0, the tokens x86_r_0 and c_0 are the first and the second operand of the token INT_SUB, respectively. Correspondingly, there are data relations between them. Consequently, we obtain the token-based semantic graph (TSG).
Step ❹. We first create virtual basic block nodes for each basic block, and then introduce dominance and post-dominance relations between them by utilizing dominance analysis. These two relations not only cover the original control flow relations but also provide more stable semantic information than control flow relations. For example, there is no directed edge between BB1 and BB4 in Fig. 3b, but in Fig. 3d, two directed edges are established between virtual basic block nodes BB1 and BB4 due to their dominance and post-dominance relations. As a result, we construct the basic block-based semantic graph (BBSG).
Step ❺. We introduce contain relation between the opcode token on the TSG and the virtual basic block node on the BBSG. A directed edge from a virtual basic block node to an opcode node if and only if the instruction of the opcode is contained within the corresponding basic block. Finally, we obtain the DESG.
In summary, the DESG includes three types of nodes and five types of edges. First, a node in a DESG can be one of the following types: (1) Virtual basic block node (e.g., BB1), where we use the index of the basic block as the node attribute. (2) Opcode node (e.g., INT_SUB), with its mnemonic serving as the node attribute. (3) Operand node (e.g., x86_r_0), where we use the token after varnode normalization strategies as the node attribute. Second, five types of edges can be defined between nodes in a DESG: (1) Data relation, which reveals the node using values defined by another node. (2) Effect relation, which imposes the necessary execution order. (3) Contain relation, indicating that an opcode node is contained within a virtual basic block node. (4) Post-dominate relation, which reveals the post-dominance relations between virtual basic block nodes. (5) Dominate relation, which indicates the dominance relations between virtual basic block nodes.
Input: A binary function $f$.
Output: A DESG $\mathcal{G}_{\text{DESG}}$ for representing the binary function $f$.
1 Let $\mathcal{G}_{\text{DESG}}$, $\mathcal{G}_{\text{ISG}}$, and $\mathcal{G}_{\text{BBSG}}$ be empty graphs;
2 Decompile $f$ into the CFG $\mathcal{G}_{\text{CFG}}$;
3 Apply varnode normalization strategies on $\mathcal{G}_{\text{CFG}}$;
4
5for $block\in\mathcal{G}_{\textnormal{CFG}}$ do
6       for $inst\in\textnormal{GetInstruction}(block)$ do
7             $\mathcal{G}_{\text{ISG}}$.addNode($inst$);
8
9            for $op\in\textnormal{the source operands of}\ inst$ do
10                   if $op$ is defined by instruction $defInst$ then
11                         $\mathcal{G}_{\text{ISG}}$.addDataEdge($inst$, $defInst$);
12                         if $defInst$ accesses memory then
13                               $\mathcal{G}_{\text{ISG}}$.addEffectEdge($inst$, $defInst$);
14
[…]
16 for $inst\in\mathcal{G}_{\textnormal{TSG}}$ do
17       Split the opcode of $inst$ as a node $opc$;
18       Refine the effect edge of $inst$ into the effect edge of $opc$;
19       Split the target operand of $inst$ as a node $op$;
20       $\mathcal{G}_{\text{TSG}}$.addDataEdge($op$, $opc$);
21       for $op\in\textnormal{the source operands of}\ inst$ do
22             Split $op$ as a node;
23             $\mathcal{G}_{\text{TSG}}$.addDataEdge($opc$, $op$);
24
[…]
26 $\mathcal{T}_{\textnormal{PD}}\leftarrow\textnormal{getPostDominatorTree}(\mathcal{G}_{\text{CFG}})$;
27 for $block\in\mathcal{G}_{\textnormal{CFG}}$ do
28       Create a virtual basic block node $bb_{i}$ for $block$;
29       $\mathcal{G}_{\text{BBSG}}$.addNode($bb_{i}$);
30
31      $bb_{j}\leftarrow\mathcal{T}_{\textnormal{D}}.\textnormal{getPredecessor}(bb_{i})$;
32       $\mathcal{G}_{\text{BBSG}}$.addDomainceEdge($bb_{j}$, $bb_{i}$);
33
34      $bb_{j}\leftarrow\mathcal{T}_{\textnormal{PD}}.\textnormal{getPredecessor}(bb_{i})$;
35
36      $\mathcal{G}_{\text{BBSG}}$.addPostDominanceEdge($bb_{j}$, $bb_{i}$);
37
38Merge $\mathcal{G}_{\text{TSG}}$ and $\mathcal{G}_{\text{BBSG}}$ into $\mathcal{G}_{\text{DESG}}$ using contain edges;
Algorithm 1 The construction algorithm of the DESG
3.3. Model Training
GGNN (Li et al., 2015) is applied to generate embedding. We first present the GGNN’s structure before discussing the model training.
3.3.1. GGNN
GNN is a neural network model for encoding nodes and edges, and learning their embeddings. GNN consists of multiple message passing layers and a graph pooling layer. GGNN is a variant of GNN, which extends GNN by incorporating a Gated Recurrent Unit (GRU) activation in the message passing phase to capture information from neighboring nodes. Following (He et al., 2024; Li et al., 2023; Guo et al., 2022), we adopt GGNN to learn the DESG’s semantics.
Message Passing. A node updates its hidden state by aggregating the hidden states of its neighbors along with its state from the previous layer. Let $\mathcal{N}$ be the set of all nodes and $\mathbf{h}_{u}^{l}$ be the hidden state of node $u$ at the $l$-th layer of GGNN. The aggregation of the neighbor’s hidden states for node $u$ is formulated as Eqn. (1).
(1) $\mathbf{m}_{u}^{l}=\sum_{v\in\text{Edge}(u)}f\left(\mathbf{h}_{u}^{l},\mathbf{h}_{v}^{l},\mathbf{e}_{v,u},\mathbf{e}_{u,v}\right)$
where $\text{Edge}(u)$ denotes the set of incoming and outgoing edges for node $u$, $\mathbf{e}_{u,v}$ represents the edge embedding of the direct edge $(u,v)$, and $f$ refers the application of two separate multilayer perceptrons used to aggregate incoming and outgoing edges, respectively. Then, in the $(l+1)$-th layer, the hidden state of node $u$ is formulated as Eqn. (2).
(2) $\mathbf{h}_{u}^{l+1}=\text{GRU}\left(\mathbf{h}_{u}^{l},\mathbf{m}_{u}^{l}\right)$
where GRU regulates information flow through gating mechanisms, retaining key features while discarding noise.
Graph Pooling. After message passing layers, the final step is to pool all node embeddings into a graph embedding. Softmax (Li et al., 2020) is a widely used pooling method, which is formulated as Eqn. (3).
(3) $\text{Softmax}(\mathcal{H}\mid\beta)=\sum_{\mathbf{h}_{u}^{L}\in\mathcal{H}}\frac{\exp\left(\beta\odot\mathbf{h}_{u}^{L}\right)}{\sum_{\mathbf{h}_{v}^{L}\in\mathcal{H}}\exp\left(\beta\odot\mathbf{h}_{v}^{L}\right)}\odot\mathbf{h}_{u}^{L}$
where $\mathcal{H}$ refers to the set of all node hidden states and $\beta$ is a parameter known as the inverse temperature. Multi-head attention is a core technique in the transformer architecture (Vaswani, 2017), allowing each head to learn different attention distributions. Similar to (He et al., 2024), we integrate softmax pooling with multi-head attention, forming a stack of multiple softmax layers. We first transform the hidden state $\mathbf{x}_{u}^{k}=\text{LayerNorm}\big{(}\text{ReLU}(\text{Linear}_{k}(\mathbf{h}_{u}^{r}))\big{)}$ using $k$ linear layers, followed by a Relu activation and layer normalization. Then, the calculation of every head is formulated as Eqn. (4).
(4) $\mathbf{e}_{g}^{k}=\text{Softmax}\left(\left\{\mathbf{x}_{u}^{k}\mid u\in\mathcal{N}\right\}\mid\beta^{k}\right)$
where $k$ denotes the index of the head. Finally, all heads are concatenated and passed through a linear layer to produce the final graph embedding $\mathbf{e}_{g}$, formulated as $\mathbf{e}_{g}=\text{Linear}\big{(}\text{Concat}(\mathbf{e}_{g}^{1},\mathbf{e}_{g}^{2},\dots,\mathbf{e}_{g}^{h})\big{)}$, where $h$ denotes the number of heads.
3.3.2. Model Training
We employ distance weighted negative sampling (Wu et al., 2017) with a margin-based pairwise loss (Li et al., 2019) to train the GGNN. This approach helps the model effectively distinguish between positive and negative samples, enhancing the capability of the embeddings. The loss function is formulated as Eqn. (5).
(5) $\min_{\theta}\mathcal{L}(\theta)=\sum_{\langle\mathbf{p},\mathbf{q}\rangle\in\mathcal{B}}\big{(}\text{positive}(\mathbf{p},\mathbf{q})+\text{negative}(\mathbf{p})\big{)}$
where $\mathcal{B}$ represents a training batch, consisting of a sequence of pairs, each in the form of $\langle\mathbf{p},\mathbf{q}\rangle$, with $\mathbf{p}$ and $\mathbf{q}$ denoting the graph embeddings originating from the same function. The positive computation is used in the loss function as $\text{positive}(\mathbf{p},\mathbf{q})=\max\big{(}1-\text{cos}(\mathbf{p},\mathbf{q})-m,0\big{)}$ and the negative sampling is applied as $\text{negative}(\mathbf{p})=\max\big{(}\text{cos}(\mathbf{p},\text{sampling}(\mathbf{p}))-m,0\big{)}$, where it samples a negative sample pair for $\mathbf{p}$ and calculates the cosine similarity, $m$ defines the margin that enforces a separation between positive and negative samples, ensuring positive pairs are closer and negative pairs are sufficiently separated.
4. Experimental Setup
4.1. Dataset
Our dataset is built on the BinKit dataset (Kim et al., 2023), a large-scale benchmark for BCSA. It includes normal and obfuscation versions generated using the same optimization levels and ISAs. Normal versions include 3 ISAs (MIPS, ARM, and x86), 9 compiler versions (5 GCC, 4 Clang versions), and 5 different optimization levels (O0 to O3 and Os). Based on Obfuscator-LLVM (O-LLVM) (Junod et al., 2015), obfuscation versions includes 4 obfuscation options: instruction substitution (SUB), bogus control flow (BCF), control flow flattening (FLA), and a combination of all obfuscation options (SUB+BCF+FLA), referred to as ALL. To evaluate the effectiveness of ORCAS in analyzing the semantic similarity of binary functions across different obfuscation options, optimization levels, and ISAs, we integrate binaries from both normal and obfuscation versions. For our experiments, we select binaries from the x86-64 and ARM-64 architectures, compiled with Clang for normal versions and O-LLVM (BCF, FLA, SUB, and ALL) for obfuscation versions, specifically with O0 and O3 optimization levels, covering 578 binaries.
4.2. Metrics
We choose several distinct metrics to comprehensively evaluate the performance of BCSA.
4.2.1. One-to-One Similarity Matching
Several works (Haq and Caballero, 2021) utilize accuracy or recall to evaluate the performance of the model in similarity matching. However, these metrics at a fixed threshold may not fully reflect a model’s ability to distinguish similar and dissimilar pairs. Instead, we evaluate and compare the similarity matching performance of different approaches using the Precision-Recall (PR) curve. Let $TP$, $FP$, and $FN$ represent the number of True Positive, False Positive, and False Negative binary function pairs inferred by the model at a given threshold, respectively, we have $\text{Precision}={TP/(TP+FP)}$ and $\text{Recall}={TP/(TP+FN)}$. To save space, we employ the the Area Under the PR Curve (PR-AUC) as a comprehensive metric to summarize the overall performance, providing a more comprehensive comparison across different approaches.
4.2.2. One-to-Many Similarity Searching
Given a pool of original function $\{f_{1},\dots,f_{n}\}$, and a pool of target functions $\{g_{1},\dots,g_{n}\}$, where one of these target functions is $f_{i}^{gt}$, corresponding to the group truth $f_{i}$. The binary function similarity searching task aims to retrieve the top-k functions most similar to $f_{i}$ from target functions, ranked by their similarity score. The performance of the task is evaluated using Recall at top-k (Recall@k) and the Mean Reciprocal Rank (MRR). Recall@k measures the model’s ability to include the group truth function within top-k results, which can be calculated by Eqn. (6).
(6) $\text{Recall}@k={1\over n}\sum_{i=1}^{n}{\mathbb{I}\big{(}{\text{Rank}(f_{i}^{gt})}\leq k\big{)}}$
where the indicator function $\mathbb{I}(x)$ maps boolean values to binary integers, with 0 for False and 1 for True. MRR evaluates the model based on the rank of its first correct answer, as specified in Eqn. (7).
(7) $\text{MRR}={1\over n}\sum_{i=1}^{n}{1\over{\text{Rank}(f_{i}^{gt})}}$
4.3. Baselines
We compare ORCAS to the following eight baselines.
SAFE (Massarelli et al., 2022) employs a recurrent neural network (RNN) augmented with attention mechanisms. It generates the embeddings of binary functions from the assembly instructions.
Gemini (Xu et al., 2017b) takes the CFG as input, assigning manually crafted features to each basic block to construct an attributed CFG. It then uses GNN to calculate the similarity between binary functions by introducing a Siamese network.
Asm2vec (Ding et al., 2019) uses random walks model the CFG as multiple sequences and applies the PV-DM model (Le and Mikolov, 2014) to learn the embeddings of binary function.
Graph Matching Networks (GMN) (Li et al., 2019) proposes a variant of the GNN that incorporates a cross-graph attention-based matching mechanism. It takes pairs of CFG as input and computes a similarity score between them.
Trex (Pei et al., 2023) first employs a hierarchical transformer-based model to capture the execution semantics of binary functions from micro traces. Then, it transfers the acquired semantic knowledge to match semantically similar binary functions. Our implementation is directly based on the source code released by the authors. [^3] https://github.com/CUMLSec/trex
HermesSim (He et al., 2024) introduces novel SOG to represent binary function because it has well-defined semantic structures. Then, it utilizes GGNN to generate the embeddings of binary functions and calculate their cosine similarity. Our implementation is based on the source code released by the authors. [^4] https://github.com/NSSL-SJTU/HermesSim
jTrans (Wang et al., 2022) proposes a jump-aware transformer-based model pre-trained on a large dataset, BinaryCorp, comprising approximately 21 million x86 binary functions. We fine-tune the pre-trained model released by the anthors on our dataset. [^5] https://github.com/vul337/jTrans
CRABS-former (Feng et al., 2024) is a BCSA model built on top of jTrans. It introduces several assembly normalization strategies for x86-64 and ARM-64 to improve the performance of BCSA. We implement CRABS-former and fine-tune it on our dataset.
Our implementation of SAFE, Gemini, Asm2vec, and GMN is based on the implementation provided by (Marcelli et al., 2022), with the default parameter settings. [^6] https://github.com/Cisco-Talos/binary_function_similarity
5. Evaluation
This section aims to address the following research questions (RQs).
  • RQ1: How performance is ORCAS on one-to-many similarity searching for binary functions compared with other baselines? (§5.1)
  • RQ2: How accurate is ORCAS on one-to-one similarity matching of obfuscated binary function compared with other baselines? (§5.2)
  • RQ3: How performance is the impact of incorporating dominance analysis on the performance of ORCAS? (§5.3)
  • RQ4: How effective is ORCAS for real-world vulnerability detection compared to other baselines? (§5.4)
We conduct all training and inference experiments on an x86 server running Ubuntu 20.04 LTS, equipped with an Intel Xeon 32-core 2.90GHz CPU, 256GB RAM, and 2 NVIDIA GeForce RTX 3090 GPUs.
5.1. RQ1: Searching against all Binary Functions
We evaluate the performance of ORCAS in searching similar binary functions across different obfuscation options, optimization levels, and ISAs for a given query function. Specifically, we evaluate ORCAS through three BCSA searching subtasks: (1) cross-ISA (XA), (2) cross-optimization level (XO), and (3) cross-ISA, optimization level, and obfuscation option (XM). Notably, obfuscated binary functions are present in all three subtasks. We conduct these evaluations on our dataset, setting the pool size for these subtasks to range from 10 to 2,000. Due to Trex’s slow sampling rate, evaluation is limited to pool sizes below 1,000. Thus, we obtain the Recall@1 and MRR for different pool size settings, as shown in Tables 3, 4, and 5.
In the XA searching subtask, ORCAS achieves 39.8%/41.7% higher Recall@1/MRR than CRABS-former when pool size is set to 2,000 via its architecture-agnostic IR design, while surpassing SOTA approach, HermesSim, by 9.6%/7.4%. Notably, Asm2vec’s lack of cross-architecture support precluded evaluation in this subtask. In the XO searching subtask, ORCAS achieves 12.8%/11.0% improvement Recall@1/MRR, when pool size is set to 2,000, compared to HermesSim. In the XM searching subtask, the Recall@1/MRR of ORCAS surpass HermesSim by 9.6%/8.4% at a pool size of 2,000.
Table 3. Results of different BCSA approaches on XA subtask.
Models Pool size = 10 Pool size = 100 Pool size = 1,000 Pool size = 2,000
Recall@1 MRR Recall@1 MRR Recall@1 MRR Recall@1 MRR
SAFE 0.506 0.685 0.133 0.269 0.023 0.069 0.011 0.042
Gemini 0.741 0.845 0.367 0.516 0.133 0.232 0.099 0.173
Asm2vec - - - - - - - -
GMN 0.788 0.879 0.326 0.509 0.069 0.168 0.040 0.108
Trex 0.455 0.594 0.193 0.305 - - - -
jTrans 0.467 0.668 0.089 0.225 0.014 0.049 0.008 0.030
CRABS-former 0.836 0.904 0.549 0.671 0.308 0.412 0.242 0.337
HermesSim 0.948 0.968 0.870 0.909 0.654 0.761 0.544 0.680
ORCAS 0.961 0.974 0.913 0.937 0.744 0.825 0.640 0.754
Table 4. Results of different BCSA approaches on XO subtask.
Models Pool size = 10 Pool size = 100 Pool size = 1,000 Pool size = 2,000
Recall@1 MRR Recall@1 MRR Recall@1 MRR Recall@1 MRR
SAFE 0.405 0.613 0.085 0.210 0.010 0.044 0.007 0.027
Gemini 0.527 0.681 0.144 0.274 0.033 0.080 0.015 0.048
Asm2vec 0.433 0.646 0.100 0.210 0.034 0.065 0.022 0.046
GMN 0.610 0.755 0.197 0.359 0.032 0.100 0.020 0.064
Trex 0.387 0.532 0.122 0.211 - - - -
jTrans 0.401 0.604 0.111 0.224 0.019 0.060 0.031 0.060
CRABS-former 0.624 0.761 0.293 0.424 0.120 0.201 0.085 0.147
HermesSim 0.900 0.934 0.788 0.842 0.579 0.677 0.476 0.595
ORCAS 0.937 0.957 0.860 0.896 0.690 0.769 0.604 0.705
Table 5. Results of different BCSA approaches on XM subtask.
Models Pool size = 10 Pool size = 100 Pool size = 1,000 Pool size = 2,000
Recall@1 MRR Recall@1 MRR Recall@1 MRR Recall@1 MRR
SAFE 0.510 0.685 0.123 0.252 0.030 0.070 0.016 0.045
Gemini 0.473 0.646 0.157 0.275 0.058 0.103 0.046 0.078
Asm2vec - - - - - - - -
GMN 0.726 0.832 0.319 0.481 0.055 0.150 0.032 0.097
Trex 0.302 0.460 0.154 0.240 - - - -
jTrans 0.277 0.511 0.036 0.128 0.003 0.019 0.001 0.013
CRABS-former 0.509 0.682 0.160 0.292 0.036 0.089 0.027 0.069
HermesSim 0.908 0.939 0.816 0.859 0.642 0.723 0.557 0.660
ORCAS 0.946 0.963 0.878 0.907 0.737 0.804 0.653 0.744
5.2. RQ2: Matching Obfuscated Binary Functions
Table 6. PR-AUC socres of matching against obfuscated binary functions on different obfuscation techniques.
Models Bogus Control Flow (BCF) Control Flow Flattening (FLA) Instruction Substitution (SUB) SUB+BCF+FLA (ALL)
Readline Sed Sharutils Tar Avg. Readline Sed Sharutils Tar Avg. Readline Sed Sharutils Tar Avg. Readline Sed Sharutils Tar Avg.
SAFE 0.072 0.079 0.095 0.078 0.081 0.080 0.094 0.075 0.074 0.080 0.142 0.142 0.153 0.153 0.147 0.054 0.059 0.055 0.046 0.053
Gemini 0.062 0.089 0.083 0.063 0.074 0.026 0.026 0.025 0.022 0.024 0.515 0.557 0.570 0.548 0.547 0.018 0.020 0.017 0.016 0.017
Asm2vec 0.013 0.021 0.028 0.027 0.022 0.011 0.012 0.013 0.010 0.011 0.292 0.422 0.371 0.499 0.396 0.007 0.007 0.007 0.007 0.007
GMN 0.329 0.282 0.359 0.301 0.317 0.165 0.180 0.229 0.184 0.189 0.513 0.488 0.604 0.529 0.533 0.164 0.137 0.184 0.092 0.144
jTrans 0.055 0.067 0.126 0.108 0.089 0.063 0.071 0.100 0.097 0.082 0.373 0.459 0.524 0.570 0.481 0.015 0.022 0.020 0.027 0.021
CRABS-former 0.040 0.070 0.057 0.088 0.063 0.015 0.016 0.012 0.028 0.017 0.620 0.686 0.755 0.772 0.708 0.014 0.014 0.011 0.011 0.012
HermesSim 0.805 0.892 0.905 0.913 0.878 0.827 0.914 0.910 0.917 0.892 0.908 0.925 0.981 0.959 0.943 0.684 0.850 0.823 0.699 0.764
ORCAS 0.840 0.935 0.965 0.946 0.921 0.881 0.967 0.971 0.939 0.939 0.947 0.957 0.993 0.969 0.966 0.792 0.935 0.933 0.800 0.865
Similar to (Ding et al., 2019; Zhang et al., [n. d.]; Pei et al., 2023; Feng et al., 2024), we evaluate ORCAS’s ability to match obfuscated binary functions with their normal version. First, we select representative projects (i.e., Readline, Sed, Sharutils, and Tar) from our dataset. For each project, we compile all functions without applying any obfuscation techniques. Next, we compile the project again using the selected obfuscation technique. Then, we can obtain both normal and obfuscated versions for each function. We link the normal and obfuscated functions to establish a one-to-one match using function symbols for evaluation. We consider function pairs compiled from the same function as positive samples, with others treated as negative samples, setting a 1:100 ratio of positive to negative samples. Finally, we use the PR-AUC score as the evaluation metric for each project.
Table 6 presents the PR-AUC scores of ORCAS and other baselines on different obfuscation techniques. Notably, ORCAS outperforms HermesSim with average PR-AUC improvements of 4.3%, 4.7%, 2.3%, and 12.1% on BCF, FLA, SUB, and ALL obfuscation respectively.
Figure 4. Average PR-AUC socres of ORCAS, ORCAS-RD on different obfuscation techniques.
5.3. RQ3: Ablation Experiment
ORCAS employs a key technique by incorporating both dominance and post-dominance relations into the DESG. To demonstrate the effectiveness, we conduct ablation experiments using ORCAS-RD, a modified version that specifically removes the dominance and post-dominance relations from the DESG while retaining all other components. The experimental setup follows the same configuration as used in RQ2.
Fig. 4 presents a comparison of PR-AUC scores between ORCAS and ORCAS-RD. The results reveal a performance degradation on different obfuscation techniques. Specifically, ORCAS-RD achieves average PR-AUC score degradations of 1.7%, 1.7%, 1.6%, and 4.0% on BCF, FLA, SUB and ALL obfuscation, respectively, compared to ORCAS. The most significant performance gap occurs on ALL obfuscation, demonstrating that dominance analysis provides the greatest benefits when dealing with multiple combined obfuscation techniques. Therefore, these results demonstrate that ORCAS’s key technique effectively improves the accuracy and robustness of similarity detection between obfuscated binary functions.
5.4. RQ4: Real-World Vulnerability Search
Previous works (Luo et al., 2023; Wang et al., 2022, 2024) utilize 1-day vulnerability detection to evaluate the effectiveness of BCSA approaches in real-world tasks. Similar to (Wang et al., 2022), we perform our evaluation on eight Common Vulnerabilities and Exposures (CVEs) from a public dataset presented in (Wang et al., 2019). We produce 10 variants for each function by O-LLVM (Junod et al., 2015) with 2 ISAs (x86-64 and ARM-64) and 5 obfuscation options (disable obfuscation, enable SUB, BCF, FLA, and ALL obfuscation options). We enable -O3 optimization option to make them closer to the release version as suggested in (Kim et al., 2023), and apply -fno-inline-functions option to facilitate locating a certain vulnerable function. Thus, we obtain a challenging obfuscated real-world vulnerability dataset, which has been released to facilitate future research on obfuscated binary code analysis. [^7] https://anonymous.4open.science/r/ORCAS-CF05
We provide details for evaluating the effectiveness of vulnerability detection. Similar to (Wang et al., 2022, 2024), for each CVE, we first extract the functions in the 10 project variants to build a function pool $P$ and mark the 10 vulnerable functions in $P$. Then, we search in $P$ using a variant of the vulnerable function and it returns the top 10 similar binary functions. The number of function pools for each CVE varies from 18,719 to 57,673. Finally, assuming $m$ of the returned binary functions are marked, we calculate the recall rate by $m/10$.
In our experiment, we select the vulnerable functions compiled with ALL obfuscation option as our queries, then use the average recall rate as the evaluation metric. Fig. 5 illustrates the results of the recall rate for each CVE. We compare ORCAS to two outstanding baselines: GMN and HermesSim. ORCAS outperforms both baseline models, e.g., for CVE-2017-14152, from the openjpeg project, which includes 18,719 functions, ORCAS achieved a 100% recall rate, meaning it can recall all variants of the vulnerable functions, improving by 43% over HermesSim and by 72% over GMN. This demonstrates that ORCAS can be effectively applied to real-world vulnerability search.
Figure 5. Recall rate of real-world vulnerability search.
6. Related Work
This section provides a brief overview of recent BCSA approaches.
6.1. Binary Code Representation in BCSA
Various representations are employed to capture the semantics of binary code, including trees, instruction streams, and graphs.
Trees. TEDEM (Pewny et al., 2014) transforms basic block instructions into equations through symbolic simplification to generate expression trees, then measures similarity using tree edit distance. Recent approach by Asteria (Yang et al., 2021) finds that abstract syntax tree (AST) exhibit greater stability than CFG cross-ISA, employing Tree-LSTM networks to encode AST of a binary function into embeddings. Building on Asteria, Asteria-Pro (Yang et al., 2023a) enhances detection efficiency by incorporating pre-filtering and re-ranking modules. RCFG2Vec (Li et al., 2024) converts each assembly instruction into a three-level syntax tree, where the root represents the instruction itself and child nodes denote its opcode or operands, subsequently generating three types of embeddings.
Instruction streams. Both assembly code and IR code are widely adopted for binary representation (Massarelli et al., 2022; Li et al., 2021; Wang et al., 2022; Ahn et al., 2022; Luo et al., 2023). Early approach SAFE (Massarelli et al., 2022) employed skip-gram models, which predict surrounding instructions from a target instruction, to generate instruction embeddings. With advances in NLP technologies, pre-trained models have demonstrated superior performance. Researchers (Li et al., 2021; Wang et al., 2022; Ahn et al., 2022; Luo et al., 2023) develop specialized pre-training tasks to enhance semantic capture, with PalmTree (Li et al., 2021) implementing both context window prediction and def-use prediction tasks. Moreover, Vulhark (Luo et al., 2023) further enhancing IR through microcode-based optimizations including root-operand analysis and instruction simplification. However, substituting complex instructions for simpler ones significantly degrades representation effectiveness under obfuscation.
Graphs. CFG serve as a prevalent representation owing to their cross-ISA ability (He et al., 2024), often extended through integration with other features. Early approaches (Feng et al., 2016; Xu et al., 2017b) construct a attributed CFG using manually designed basic block features, while recent approaches integrate CFG with CG and DFG (Kim et al., 2022; Duan et al., 2020; Guo et al., 2022). DeepBinDiff (Duan et al., 2020) merges CG and CFG to construct a inter-procedural CFG. XBA (Kim et al., 2022) propose a binary disassembly graph that leverages basic blocks and external functions. Additionally, HermesSim (He et al., 2024) proposes a SOG that decomposes instructions into operand and opcode nodes to reveal their relations. These approaches achieve promising performance but become unstable under obfuscation due to their reliance on original control flow.
6.2. Resilience to Obfuscation in BCSA
Recent approaches have been addressing the challenges posed by obfuscation. Asm2vec (Ding et al., 2019) applies random walk and edge coverage on the CFG to sample instruction sequences, then trains PV-DM model to generate function embeddings. CoP (Luo et al., 2017) integrates program semantics with fuzzy matching of the longest common subsequence to capture basic block semantics. BinFinder (Qasem et al., 2023) first identifies a set of engineered interpretable features, then trains a neural network model using a siamese network architecture. GTrans (Zhang et al., [n. d.]) leverages the D-Tree structure and employs three encoding strategies to train a graph attention network. HermesSim (He et al., 2024) introduces a SOG to represent binary functions and generates function embeddings using GGNN. Our experimental results shows its resilience against obfuscation. Sem2vec (Wang et al., 2023) learns continuous and short execution traces from the CFG, extracts constraints through symbolic execution. Trex (Pei et al., 2023) learns program execution semantics from a mixture of regular and forced-execution traces, and then transfers the acquired semantic knowledge to match semantically similar binary functions. Although these approaches are resilient to obfuscation and can detect similarities between obfuscated binary, they still fall short of capturing the complete semantics of obfuscated binary, limiting their effectiveness when multiple obfuscation techniques are applied simultaneously.
7. Conclusion
This paper proposes ORCAS, a novel obfuscation-resilient BCSA model based on Dominance Enhanced Semantic Graph (DESG). The DESG captures binaries’ implicit semantics, including inter-basic block, inter-instructions and instruction-basic block relations, then gated graph neural networks are utilized to generate the embedding. We contribute the obfuscated real-world vulnerability dataset for evaluating 1-day vulnerability detection on different obfuscation techniques. Extensive comparative experiments and ablation experiment demonstrate ORCAS’s performance, showing that ORCAS outperforms baselines on both the BinKit dataset and our original obfuscated real-world vulnerability dataset.